OpenStudy (anonymous):

Graph each equation by plotting ordered pairs. 30. y = x + 2

OpenStudy (johnweldon1993):

so lets just plug in numbers for x 0 1 2 3 y = x + 2 = 0 + 2 ....so at x = 0, y = 2 = 1 + 2....so at x = 1, y = 3 = 2 + 2....so at x = 2, y = 4 =3 + 2.....so at x = 3, y = 5 so on an so forth.....then connect the points :)

OpenStudy (anonymous):

thank you :) but how did you get = 0 + 2?

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (anonymous):

all he did was just put in differant numbers for x... then you can plt those points adn then connect them and get a line. ;)

OpenStudy (johnweldon1993):

yeah sorry about that..... well 0 + 2....that was from when i made x = 0 your equation was y = x + 2 right? so when x = 0...you have y = 0 + 2 better?

OpenStudy (johnweldon1993):

@foreverkamixo

OpenStudy (anonymous):

yeah :) thanks!

OpenStudy (johnweldon1993):

anytime :)